Find the LCM for the pair of polynomial: 7m-21 and 14m-42.
Answer by venugopalramana(3286)

Find the LCM for the pair of polynomial: 7m-21 =
=7(M-3)
and 14m-42 =14(M-3)
(M-3) IS A COMMON FACTOR AND LCM OF 7 AND 14 IS 14...HENCE LCM IS
14(M-3)